The units digit of a two digit number is 1 less than twice the tens digit. The square of the tens digit is 9 more than the units digit. Find the number.
Answer by jorel1380(3719)

2x=y+1
x2=y+9
2x-1=y
x2=2x-1+9
x2-2x-8=0
(x-4)(x+2)=0
x=4 or -2
Throwing out the negative number, we get x=4, y=7; thus the number is 47..